OpenStudy (anonymous):

solve the equation and chec the solution s-27=-40

jimthompson5910 (jim_thompson5910):

You need to undo that -27 so you can isolate s.

jimthompson5910 (jim_thompson5910):

So how do you undo that operation?

OpenStudy (anonymous):

S=-13 I don't really know how to explain how I got it... :(

jimthompson5910 (jim_thompson5910):

You undo it by adding 27 to both sides You want the left side to have -27+27 = 0 The 0 will go away (leaving just s behind)

jimthompson5910 (jim_thompson5910):

s-27=-40 s-27+27=-40+27 s + 0 = -13 s = -13

OpenStudy (anonymous):

Yea.. I just know what it is.. :)